# Improved Patch Denoising Diffusion Probabilistic Models for Magnetic Resonance Fingerprinting

### v1

@@ 4.2 Tested Algorithms
We compare our approach with compressed sensing and deep learning methods for MRF reconstruction. Tested methods used SVD dimensionality reduction with $s=5$ to reconstruct TSMIs (see 2.1), except for SCQ which uses a different dimensionality reduction learned by a neural network. Below we describe each tested baselines:
@@ 4.2.1 MRF-IDDPM
Our implementations used the publicly available code for IDDPM [^1] https://github.com/openai/improved-diffusion in PyTorch. The architecture follows the U-Net structure described in [21], consisting of four downsampling/upsampling blocks (starting with 128 channels), two residual blocks, and one attention head introduced at a resolution of $16\times 16$. During inference, $K=50$ IDDPM sampling steps were used instead of the $T=1000$ diffusion steps. Our base model reconstruct the TSMI from overlapping patches of size 64 and stride 8. To evaluate performance, we run the reconstruction process ten times with different random initializations of $\mathbf{x}_{T}$, and averaged the TSMI results before dictionary matching. Our ablation studies (section 4.5) demonstrate that these settings are sufficient to reconstruct high-quality tissue parameter maps.
@@ 4.2.3 LRTV
[40] A compressed sensing method based on convex optimisation to reconstruct TSMIs from undersampled k-space measurements by enforcing k-space data consistency and applying spatial Total Variation regularization.
@@ 4.2.4 TSMI restoration CNN
We employed Zhang’s Deep Residual U-Net (DRUNet) model [42], a state-of-the-art CNN recently used for MRF reconstruction in [16, 43]. We trained the network for TSMI restoration task; similar to our MRF-IDDPM, the inputs are aliased-contaminated TSMIs from gridding reconstruction of faster MRF acquisitions, and outputs are the reference TSMIs from longer acquisitions. Similar to [43], the residual U-Net architecture had four down/upsampling blocks staring with 32 feature channels, ReLU activations, strided convolutional layers for downsampling and transpose convolutional layers for up sampling. Model was trained using Mean Squared Error (MSE) loss and ADAM optimiser for 5,000 epochs, with tuned hyperparamters for learning rate $10^{-4}$ and batch size 32, and using the same patch-wise data augmentation strategy employed in MRF-IDDPM training.
@@ 4.2.5 SCQ
In contrast to the previous methods, the SCQ technique [13] takes the full length of the accelerated scans ($l=200$) as input and directly outputs the parameter maps of interest. Instead of using SVD, SCQ firstly compresses the TSMIs from $l$ to 46 with a fully connected neural network. This is then processed by two U-Nets to estimate the T1 and T2 tissue parameters, respectively. To implement the SCQ method we followed the description provided in [13]. Similar to our approach, SCQ models were trained on patches of size $64\times 64$, while the hyperparameters learning rate, loss function and number of epochs were grid-searched for optimal performance. We found that a learning rate of 0.01, with MSE optimisation over 5,000 epochs, yielded the best results for our data.
Method MAPE (%) RMSE (ms) NRMSE SSIM
T1 T2 T1 T2 T1 T2 TSMI T1 T2 TSMI
SCQ 9.58 23.67 16.22 9.16 0.1143 0.4354 0.9750 0.9674
MRF-IDDPM (Ours) 7.20 17.69 14.13 7.67 0.1002 0.3687 0.4142 0.9758 0.9711 0.8729
Table 1: Reconstruction metrics for TSMIs (where applicable) and tissue parameter maps (T1/T2) for the evaluated algorithms. The evaluation dataset consisted of 24 axial brain slices from two different subjects.
@@ 4.3 Evaluation metrics
We evaluate the performance of tested algorithms for T1 and T2 maps, and the TSMI reconstructions. TSMI metrics are not available for SCQ, as this method directly outputs parameter maps. For evaluations, we used two types of masks: 1) head masks, including brain and surrounding non-brain/skull tissues (excluding background), for qualitative demonstrations (Figures 1 and 2), and 2) skull-stripped (brain-only) masks for deriving quantitative metrics in Table 1. Three of the reconstruction metrics in this table relied on the error maps $\boldsymbol{\varepsilon}\in\mathbb{R}^{n\times s},{\boldsymbol{\varepsilon}_{i,j}}={\boldsymbol{\Phi}}_{i,j}-\hat{\boldsymbol{\Phi}}_{i,j}$ between the reference $\boldsymbol{\Phi}_{i,j}$ and the predicted $\hat{\boldsymbol{\Phi}}_{i,j}$ values for the voxel $j$ and channel $i$. Note that $s=5$ for TSMI reconstructions, while $s=1$ for each parameter map. For clarity, we drop the index $i$ in the T1/T2 metrics. The Mean Absolute Percentage Error (MAPE), reported only for T1 and T2 parameter maps, is defined as:
$MAPE=\frac{100}{\text{size}(\mathcal{M})}\sum_{j\in\mathcal{M}}\left|\frac{{\boldsymbol{\varepsilon}_{j}}}{{\boldsymbol{\Phi}}_{j}}\right|,$ (7)
where $\mathcal{M}$ represents the spatial mask. The Root Mean Square Error (RMSE), reported only for T1 and T2 parameter maps, defined as:
$RMSE=\left(\frac{1}{N}\sum_{j\in\mathcal{M}}{|\boldsymbol{\varepsilon}}_{j}|^{2}\right)^{1/2};$ (8)
the Normalized Root Mean Squared Error (NRMSE), reported for T1, T2 parameter maps and TSMI reconstructions, defined as:
$NRMSE=\frac{1}{s}\sum_{i=1}^{s}\frac{\|{\boldsymbol{\varepsilon}}_{i,j\in\mathcal{M}}\|_{2}}{\|{\boldsymbol{\Phi}}_{i,j\in\mathcal{M}}\|_{2}};$ (9)
and the Mean Structural Similarity Index (SSIM), reported for T1, T2 parameter maps and TSMI reconstructions. To deal with the multi-channel complex nature of the TSMI data, we compute the SSIM for the imaginary and real components for each channel separately, and report their averages.
@@ 4.4 Results and Discussions
Table 1 quantitatively compares the performance of MRF-IDDPM against other reconstruction baselines. To compliment the quantitative results, Figure 1 presents the reconstructed T1 and T2 maps from MRF-IDDPM alongside baseline methods (SVDMRF is excluded due to its less competitive performance). Figure 2 illustrates the reconstructed TSMI across $s=5$ channels comparing MRF-IDDPM results to reference images and baseline methods (with SCQ excluded, as it does not reconstruct TSMI).
@@
Visual inspection of the reconstructed maps (Figure 1) indicates seemingly good performance for T1 estimations, with errors becoming more apparent with the aid of the error maps. However artifacts are much more noticeable in the T2 reconstructions. Despite their sharpness, the reference T2 maps exhibit some high-frequency artifacts, as also observed in the reference TSMIs in Figure 2. This occurs because the MRF data ($l=1000$) used for generating the reference TSMIs and Q-Maps was also undersampled across the k-space; in MRF spiral sampling, the high-frequency regions are not densely sampled compared to the central k-space. In LRTV, the Total Variation penalty was applied mildly during reference image reconstruction to avoid over-smoothing anatomies, thereby balancing sharpness and artifacts. In contrast, IDDPM effectively removes artifacts while incurring a smaller trade-off in sharpness, even with 5x accelerated scans. Note that the LRTV or other baselines underperform MRF-IDDPM in this acceleration regime. A similar pattern is observed in the TSMI reconstructions (Figure 2), where the last two channels, arguably the weakest and most difficult to restore, are better reconstructed by MRF-IDDPM than the baselines. Overall, the MRF-IDDPM implementations provide the cleanest reconstructions among all techniques.
Refer to caption Refer to caption
Fig. 1: Reconstructed T1(left), T2(right) maps by our method and baselines along with percentage error maps for a representative brain slice from the evaluation set.
Fig. 2: Magnitudes of the reference and reconstructed TSMIs (SVD compressed to $s=5$) from a representative slice in the evaluation dataset, using different methods. Reconstructions are to basis of TSMIs of length $l=200$. These approaches aim to restore the corrupted input (condition image).
@@ 4.5 Ablation Studies
There are numerous parameters and options available for both training and sampling with diffusion models, and their selection is highly dependent on the specific application and task. In the following sections, we will discuss the most crucial parameters that were employed in our experiments. The baseline method, as described in Section 3.2, serves as the reference point for the following ablation experiments. These studies will examine the impact of modifying individual variables from the base setup. Metrics report the average performance on test dataset. Unless otherwise stated, samples were reconstructed from a single realization (i.e. reverse process is run for only one $\mathbf{x}_{T}$).
4.6 Effect of patch size
The key aspect of our approach is the use of patch-based training and sampling. To substantiate this, this study explored the optimal patch-size for best reconstructions on our data. We evaluated three input sizes: $32\times 32$, $64\times 64$ and $128\times 128$. Three MRF-IDDPM models were trained, all following the same specifications outlined in Section 3.2. The results of this study are summarised in Table 2. We observe that models trained on larger patch sizes can produce more accurate reconstructions, as they capture long-range pixel dependencies across broader spatial regions. However, training and sampling for these models require significantly more time, making them less ideal. In this study, we found that using $64\times 64$ patches provides a favourable balance between efficient training time and high reconstruction quality, with a difference of just $\sim 1\%$ in the T1/T2 MAPE compared to the model trained with a patch size of $128\times 128$.
Patch Time (s) MAPE (%) NRMSE
@@ 4.6 Effect of patch size
Table 2: Running times and reconstruction errors for TSMIs and T1/T2 parameter maps from baseline diffusion models trained on different spatial patch dimensions. Training time (T) is to complete 1k iterations, inference time (T) is to reconstruct one image. Metrics are averaged over test dataset.
4.7 Effect of patch stride size
@@
In our approach, the final output is assembled from individual reconstructed patches, with quality determined by both patch size and stride. Selecting the degree of overlap, i.e., the patch stride, requires balancing image quality with reconstruction time. High overlap improves reconstruction quality but can drastically increase the runtime. Conversely, using low to no overlap may speed up sampling but often leads to blocking image artifacts and degraded reconstruction performance. Therefore, an optimal stride size is necessary to achieve high-quality outputs while maintaining computational efficiency. We explored this trade-off by varying the stride size during IDDPM image reconstruction from 2 to 64 (non-overlapping). Figure 3 displays the reconstruction times alongside the performance metrics for T1, T2, and TSMI for each tested stride size. Both MAPE and SSIM metrics indicate a good trade-off between reconstruction accuracy vs. computation time for a stride size of 8. Stride sizes larger than 8 made reconstruction quality decline. Smaller stride sizes did not much improve reconstruction while had increased the computation time more than an order of magnitude.
@@
Fig. 3: Effect of stride size on reconstruction time (black curve), as well as the reconstruction NMRSE (left) and SSIM (right) metrics for TSMI (bottom) and and averaged T1 and T2 maps (top, labeled as (T1+T2)/2). All metrics are averaged over the testing dataset.
4.8 Number of time steps
@@
Fig. 4: Effect of timesteps (K) on NRMSE (left) and SSIM (right) on the T1/T2 reconstruction maps (top) and the enhanced TSMI (bottom). The metrics are the average across the testing dataset.
@@
We also examined the number of IDDPM time steps $K$ required for sampling accurate MRF reconstructions. The results are presented in Figure 4, showing NRMSE and SSIM as a function of $K$ for the estimations of T1/T2 maps and the TSMIs. As expected, the time required for a complete pass of the reverse process for a single image increases linearly with $K$. Reconstruction NRMSEs rapidly drop with only a few sampling steps, plateauing promptly after this fall. Similarly, the SSIM metrics improve rapidly, although it plateaus later than the NRMSE. The plots suggest that at least 30 sampling steps are enough for the reverse process. We used $K=50$ in all other experiments throughout, as a common choice in the literature [29, 44] while still achieving a good trade-off between the reconstruction accuracy and computation time.
4.9 Estimation of missing timeframes
Given the same degraded TSMI inputs of length $l=200$ from the gridding/adjoint operator, the performance of our MRF-IDDPM model and DRUNet was compared on two TSMI reconstruction tasks, each one targetting high-quality SVD compressed TSMI of 1) the same length as input, i.e. $l=200$; and 2) compressed from longer acquisition with $l=1000$. This implies that there are two different basis, $\textbf{V}_{1000}$ and $\textbf{V}_{200}$, required to compress the dictionary $\underline{\textbf{D}}$, which is then used in DM to estimate the T1/T2 parameter maps. The results of these experiments are reported in Table 3.
Method Task MAPE (%) NRMSE

### v2

@@ 4.2 Tested Algorithms
We compare our approach with classical compressed sensing and deep learning methods for MRF reconstruction. Tested algorithms, except SCQ, used dictionary matching to map the reconstructed TSMIs to T1 and T2 maps. For this purpose, an MRF dictionary consisting of 95K fingerprints was simulated on a logarithmically-spaced grid of $(T1,T2)\in[0.01,6]\text{s}\times[0.004,4]\text{s}$) using the extended-phase-graph formalism [48]. Tested methods also used SVD dimensionality reduction with $s=5$ to reconstruct TSMIs (see 2.1), except for SCQ which uses a different dimensionality reduction learned by a neural network. Below we describe each tested baselines:
@@ 4.2.1 MRF-IDDPM
Our implementations used the publicly available code for IDDPM [^2] https://github.com/openai/improved-diffusion in PyTorch. The architecture follows the U-Net structure described in [26], consisting of four downsampling/upsampling blocks (starting with 128 channels), two residual blocks, and one attention head introduced at a resolution of $16\times 16$. During inference, $K=50$ IDDPM sampling steps were used instead of the $T=1000$ diffusion steps for training. Our base model reconstruct the TSMI from overlapping patches of size 64 and stride 8. To evaluate performance, we run the reconstruction process ten times with different random initializations of $\mathbf{x}_{T}$, and averaged the TSMI results before dictionary matching. Our ablation studies (section 4.5) demonstrate that these settings are sufficient to reconstruct high-quality tissue parameter maps.
@@ 4.2.3 LRTV
A compressed sensing method [47] based on convex optimisation to reconstruct TSMIs from undersampled k-space measurements by enforcing k-space data consistency and applying spatial Total Variation regularization.
@@ 4.2.4 TSMI restoration CNN
We employed Zhang’s Deep Residual U-Net (DRUNet) model [49], a state-of-the-art CNN recently used for MRF reconstruction in [21, 50]. We trained the network for TSMI restoration task; similar to our MRF-IDDPM, the inputs are aliasing-contaminated TSMIs from gridding reconstruction of faster MRF acquisitions, and outputs are the reference TSMIs from longer acquisitions. Similar to [50], the residual U-Net architecture had four down/upsampling blocks staring with 32 feature channels, ReLU activations, strided convolutional layers for downsampling and transpose convolutional layers for up sampling. The model was trained using Mean Squared Error (MSE) loss and ADAM optimiser for 5,000 epochs, with tuned hyperparameters for the learning rate of $10^{-4}$ and batch size 32, and using the same patch-wise data augmentation strategy employed in MRF-IDDPM training.
@@ 4.2.5 SCQ
In contrast to the previous methods, the SCQ technique [19] takes the full length of the accelerated scans ($l=200$) as input and directly outputs the parameter maps of interest. Instead of using SVD, SCQ firstly compresses the TSMIs from $l$ to 46 with a fully connected neural network. This is then processed by two U-Nets to estimate the T1 and T2 tissue parameters, respectively. To implement the SCQ we followed the description provided in [19]. Similar to our approach, SCQ models were trained on patches of size $64\times 64$, while the learning rate, loss function and number of epochs were grid-searched for optimal performance. We found that a learning rate of 0.01, with MSE optimisation over 5,000 epochs, yields the best results for our data.
@@ 4.3 Evaluation metrics
We evaluate the performance of tested algorithms for T1 and T2 maps, and the TSMI reconstructions. TSMI metrics are not available for SCQ, as this method directly outputs parameter maps. For evaluations, we used two types of masks: i) head masks, including brain and surrounding non-brain/skull tissues (excluding background) for qualitative demonstrations in Figures 1, 2 and 5, and ii) skull-stripped (brain-only) masks for deriving quantitative metrics in Table 1. Some reconstruction metrics rely on error maps $\boldsymbol{\varepsilon}\in\mathbb{R}^{n\times s},{\boldsymbol{\varepsilon}_{v,i}}={\boldsymbol{\Phi}}_{v,i}-\hat{\boldsymbol{\Phi}}_{v,i}$ between the reference $\boldsymbol{\Phi}_{v,i}$ and the predicted $\hat{\boldsymbol{\Phi}}_{v,i}$ values for the voxel $v$ and channel $i$. Note that channel $s=5$ for TSMI reconstructions, while $s=1$ for each parameter map. For clarity, we drop the index $i$ in the T1/T2 metrics. To evaluate T1 and T2 mapping performances we report the Mean Absolute Percentage Errors:
$\text{MAPE}=\frac{100}{M}\sum_{v\in\mathcal{M}}\left|\frac{{\boldsymbol{\varepsilon}_{v}}}{{\boldsymbol{\Phi}}_{v}}\right|,$ (7)
and Root Mean Square Errors $\text{RMSE}=\sqrt{\frac{1}{M}\sum_{v\in\mathcal{M}}{|\boldsymbol{\varepsilon}}_{j}|^{2}}$, where $\mathcal{M}$ represents the spatial mask and $M={\text{size}(\mathcal{M})}$. We also report the Normalized Root Mean Squared Errors $\text{NRMSE}=\frac{1}{s}\sum_{i=1}^{s}{\|{\boldsymbol{\varepsilon}}_{v,i\in\mathcal{M}}\|_{2}}/{\|{\boldsymbol{\Phi}}_{v,i\in\mathcal{M}}\|_{2}}$, and the Mean Structural Similarity Index (SSIM) for T1 and T2 mappings, and the TSMI reconstructions. To deal with the multi-channel complex nature of the TSMI data, we compute the SSIM for the imaginary and real components for each channel separately, and report their averages.
LRTV          SCQ         DRUNetMRF-IDDPMReference LRTV          SCQ          DRUNetMRF-IDDPMReference
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Fig. 1: Reconstructed T1 (left panel) and T2 (right panel) maps using different methods (columns) for three representative brain slices from the evaluation set. Every three rows correspond to one brain slice, displaying (from top to bottom): the full reconstructed map, a zoomed-in view of a region of interest (green box), and the corresponding percentage error map. Electronic zooming is recommended for detailed inspection.
Fig. 2: Magnitudes of the reference and reconstructed TSMIs (SVD compressed to $s=5$) from one representative slice in the evaluation dataset, using different methods. Methods aim to restore the corrupted input (condition image) from a gridding reconstruction.
@@ 4.4 Results and Discussions
Table 1 quantitatively compares the performance of MRF-IDDPM against other reconstruction baselines. To complement the quantitative results, Figure 1 presents the reconstructed T1 and T2 maps from MRF-IDDPM alongside baseline methods (SVDMRF is excluded due to its less competitive performance). Figure 2 illustrates the reconstructed TSMI across $s=5$ channels comparing MRF-IDDPM results to reference images and baseline methods (with SCQ excluded, as it does not reconstruct TSMI).
@@
Visual inspection of the reconstructed maps (Figure 1) indicates seemingly good performance for T1 estimations, with errors becoming more apparent with the aid of the error maps. However artifacts are much more noticeable in the T2 reconstructions. Despite their sharpness, the reference T2 maps exhibit some high-frequency artifacts, as also observed in the reference TSMIs in Figure 2. This occurs because the MRF data ($l=1000$) used for generating the reference TSMIs and Q-Maps was also undersampled across the k-space and for reference image reconstruction we applied a very mild Total Variation penalty to avoid over-smoothing the anatomies, thereby balancing sharpness and artefacts. In contrast, IDDPM effectively removes artifacts while incurring a smaller trade-off in sharpness, even with 5x accelerated scans. Note that the LRTV or other baselines underperform MRF-IDDPM in this acceleration regime. A similar pattern is observed in the TSMI reconstructions (Figure 2), where the last two channels, arguably the weakest and most difficult to restore, are better reconstructed by MRF-IDDPM than the baselines. Overall, the MRF-IDDPM implementations provide the cleanest reconstructions among all techniques.
@@ 4.5 Ablation Studies
Diffusion models offer various configuration choices for both training and sampling, with their optimal selection highly dependent on the specific application and task. In the following sections, we outline the key parameters that were employed in our experiments. The baseline setup, as described in Section 3.2, serves as the reference for the following ablation experiments, where we assess the effect of altering individual parameters from the base configuration. Metrics report the average performance on test dataset. Unless otherwise stated, here samples were reconstructed from a single realization (i.e. the reverse process is run for only one $\mathbf{x}_{T}$).
Patch Time (s) $\downarrow$ MAPE (%) $\downarrow$ NRMSE $\downarrow$
@@
Table 2: The runtimes and the T1, T2 and TSMI reconstruction errors of MRF-IDDPM models trained on different spatial patch dimensions. Training time (T) is to complete 1K iterations, inference time (I) is to reconstruct one image. Metrics are averaged over test dataset.
4.5.1 Effect of patch size
A key aspect of our approach is the use of patch-based training and sampling. In this study, we investigated the impact of different patch sizes on the performance of our method. Three input sizes were considered: $p=32\times 32$, $64\times 64$ and $128\times 128$. Accordingly, three MRF-IDDPM models were trained, all following the same specifications as in Section 3.2. The results of this study are summarised in Table 2. We observe that models trained on larger patch sizes can produce more accurate reconstructions, as they capture long-range pixel dependencies across broader spatial regions. However, training and sampling for these models require significantly more time, making them less ideal. In this study, we found that using $64\times 64$ patches provides a favourable balance between efficient training time and high reconstruction quality, with a difference of just $\sim 1\%$ in the T1/T2 MAPE compared to the model trained with a patch size of $128\times 128$.
4.5.2 Effect of patch stride size
@@
In our approach, the final output is assembled from individual reconstructed patches, with quality determined by both patch size and stride. Selecting the degree of overlap, i.e., the patch stride, requires balancing image quality with reconstruction time. High overlap improves reconstruction quality but can drastically increase the runtime. Conversely, using low to no overlap may speed up sampling but often leads to blocking image artifacts and degraded reconstruction performance. Therefore, an optimal stride size is necessary to achieve high-quality outputs while maintaining computational efficiency. We explored this trade-off by varying the stride size during IDDPM image reconstruction from 2 to 64 (non-overlapping). Figure 3 displays the reconstruction times alongside the performance metrics for T1, T2, and TSMI for each tested stride size. Both MAPE and SSIM metrics indicate a good trade-off between reconstruction accuracy vs. computation time for a stride size of 8. Stride sizes larger than 8 made reconstruction quality decline. Smaller stride sizes did not improve the reconstruction much, while increasing the computation time more than one order of magnitude.
@@
Fig. 3: Effect of stride size on the reconstruction time (black curve), as well as the reconstruction NMRSE (left) and SSIM (right) metrics for TSMI (bottom) and averaged T1 and T2 maps (top, labelled as (T1+T2)/2). Metrics are averaged over the test dataset.
@@
Fig. 4: Effect of timesteps (K) on the reconstruction time (black curve), as well as the reconstruction NMRSE (left) and SSIM (right) metrics for TSMI (bottom) and averaged T1 and T2 maps (top). Metrics are averaged over the test dataset.
Fig. 5: Average, standard deviation (STD), and absolute error maps for T1 (left panel) and T2 (right panel) reconstructions of three representative brain slices in the evaluation set. Each row corresponds to a different brain slice. The MRF-IDDPM algorithm was used to generate 10 reconstructions per slice, each initiated with a different random noise $\mathbf{x}_{T}$. Average and STD maps show the pixel-wise mean and standard deviation across these 10 reconstructions. Absolute error maps display the absolute difference between the average reconstruction and the ground truth reference from Figure 1.
4.5.3 Number of time steps
@@
We also examined the number of IDDPM time steps $K$ required for sampling accurate MRF reconstructions. Results are presented in Figure 4, showing NRMSE and SSIM as a function of $K$ for the estimations of T1/T2 maps and the TSMIs. As expected, the time required for a complete pass of the reverse process for a single image increases linearly with $K$. Reconstruction NRMSEs drop rapidly with only a few sampling steps, plateauing promptly after this fall. Similarly, the SSIM metrics improve rapidly, although they plateau later than the NRMSE. The plots suggest that 30 sampling steps are enough for the reverse process. We used $K=50$ in all other experiments throughout, as a common choice in the literature [34, 51] while still achieving a good trade-off between the reconstruction accuracy and runtime.
4.5.4 Estimation of missing time frames
So far, our base models assume that the input $\mathbf{x}_{c}$ and target $\mathbf{x}_{0}$ TSMIs have the same truncated time length ($l=200$). To expand on this, we also trained models (with the same input) to reconstruct outputs $\hat{\mathbf{x}}_{0}$ of length $l=1000$, matching the longer reference acquisition sequence. These two configurations are referred to as Task 1 ($l=200$) and Task 2 ($l=1000$). Outputs of both tasks used SVD dimensionality reduction but with different bases, $\textbf{V}_{200}$ and $\textbf{V}_{1000}$, corresponding to dictionaries with truncated and untruncated fingerprints, respectively. For each task, we trained both MRF-IDDPM and DRUNet models and compared their test performances in Table 3.
Method Task MAPE (%) $\downarrow$ NRMSE $\downarrow$